# Projection-free computation of robust controllable sets with constrained zonotopes

### v1

@@ 5.2 Convex polytopic $\mathcal{X}_{t}$
where (45c) combines (43c) and (43d) into a single step using (7c). Given representation complexities $\mathscr{C}(\mathcal{X})=(n,N_{X},M_{X})$ and $\mathscr{C}(\mathcal{U})=(m,N_{U},M_{U})$, the representation complexity of $\mathcal{K}_{t}$ grows by $(0,N_{U}+N_{X},M_{U}+M_{X}+n)$ at each step of the set recursion (see (7c) and Table 2). Consequently, for $\mathcal{G}$ with $\mathscr{C}(\mathcal{G})=(n,N_{G},M_{G})$, the set representation complexity of the inner-approximation of the $T$-step RC step is
$\mathscr{C}(\mathcal{K}^{-})=(n,\ N_{G}+T(N_{U}+N_{X}),\ M_{G}+T(M_{U}+M_{X}+n)).$ (46)
Similarly to Section 5.1, the representation complexity of $\mathcal{K}^{-}$ in (46) also grows linearly with $T$ and does not depend on the disturbance set $\mathcal{W}$.
@@
We can also use the set recursions discussed in Section 5.1 and 5.2 in conjunction with Algorithm 3 to obtain an outer-approximation to the $T$-step RC set.
@@ 6 Case studies
We now demonstrate the computational efficiency, scalability, and utility of our approach in several case studies. First we consider two case studies involving low-dimensional systems to illustrate the advantage of the proposed approach in computing time when compared to existing inner-approximating approaches based on constrained zonotopes [9], and exact approaches based on polyhedra [4]. Then, we discuss the scalability of the approach by computing the RC set for a chain of mass-spring-damper system, and empirically demonstrate that the proposed approach is numerically stable and can compute the RC sets for high dimensional systems and long horizons.
We perform the presented computations in a standard computer with Intel CPU i9-12900KF processor (3.2 GHz, 16 cores) and 64 GB RAM, running MATLAB 2022b. We use YALMIP [26] to set up the optimization problems with MOSEK [27] and GUROBI [28] as solvers. For two-dimensional plots of constrained zonotopes $\mathcal{C}$, we compute the appropriate convex polytopic approximations via support function and support vector computations (4) in $100$ equi-spaced directions in $\mathbb{R}^{2}$. We estimate the volume of the sets via sampling.
6.1 Double integrator example with convex polytopic $\mathcal{X}$
We consider the computation of RC set for a double integrator system with convex polytopic state constraints $\mathcal{X}$ and an ellipsoidal disturbance set $\mathcal{W}$. The linear time-invariant system matrices are
@@ 6.1 Double integrator example with convex polytopic $\mathcal{X}$
with the sampling period $\Delta T=0.1$, the input set $\mathcal{U}=\llbracket-2,2\rrbracket$ that is an interval, the disturbance set $\mathcal{W}=(0.1I_{2},[0;0])$ or $\mathcal{W}=\text{diag}([0.2,0.04]),[0.1;0.1])\}$ that is a circle or an ellipsoid respectively, and the state constraints $\mathcal{X}=\mathcal{G}=\llbracket-2,2\rrbracket\times\llbracket-3,3\rrbracket$ that are time-invariant, axis-aligned rectangles.
@@
We generate inner-approximations of the $T$-step RC set using the recursion in Section 5.2 for $T=20$ with an exact ellipsoidal representation of $\mathcal{W}$ and a zonotopic outer-approximation $\mathcal{W}^{+}=(G_{W},c_{W})\supset\mathcal{W}$ where the Pontryagin difference is inner-approximated using Algorithm 1. We also compare the computed sets with the exact sets computed using MPT3 where the vertex-facet enumeration was accomplished using Fourier-Motzkin elimination [4], and the inner-approximations of the RC sets using the two-stage approach with the zonotopic $\mathcal{W}^{+}$ [9]. We also compute an outer-approximation of the RC sets using Algorithm 3, as discussed in Remark 1.
@@
Table 3 shows that the proposed inner-approximation approach with ellipsoidal $\mathcal{W}$ is about two orders of magnitude faster than the exact approach [4] and the two-stage approach [9], while providing reasonably accurate inner-approximations (about $97\%$ and $77\%$ of the area of the exact RC set in the first and second case respectively). The proposed inner-approximation approach with zonotopic $\mathcal{W}^{+}$ is slightly faster than with ellipsoidal $\mathcal{W}$, but is more conservative due to the zonotopic outer-approximation of the disturbance set. The shorter computation times of the proposed approach are a direct result of Theorems 1 and 2, since the implementation of the set recursion in Section 5.2 can be accomplished in closed-form, i.e., optimization-free. The proposed outer-approximation approach is slower than the inner-approximation approach, primarily due to the use of linear programming to produce a minimal representation of the convex polyhedron in Step 3 of Algorithm 3 to manage the representation complexity. For most safe constrained control problems, an inner-approximation of the $T$-step RC set is sufficient.
@@
Figure 1 shows that the RC sets and their corresponding inner-approximations, associated with a ball-shaped $\mathcal{W}$ (left) and an ellipsoidal $\mathcal{W}$ (right). As expected, the inner-approximations of the RC set constructed using zonotopic $\mathcal{W}^{+}$ are more conservative than their ellipsoid-based counterparts in both cases. On the other hand, the proposed inner-approximations with zonotopic $\mathcal{W}^{+}$ are identical (left) or similar (right) to the inner-approximations produced by the existing two-stage approach [9], while requiring significantly shorter computation time (see Table 3).
Property Exact $\mathcal{W}$ represented as an ellipsoid Zonotopic $\mathcal{W}^{+}\supset\mathcal{W}$
RC set [4] Proposed $\mathcal{M}^{-}$ Proposed $\mathcal{M}^{+}$ Proposed $\mathcal{M}^{-}$ Two-stage [9]
True $\mathcal{W}$ is a ball (left)
Ratio of area 1 0.97 1.67 0.70 0.67
Compute time (s) 2.044 0.012 4.064 0.010 1.594
Ratio of time 177.39 1 352.66 0.86 138.29
True $\mathcal{W}$ is an ellipsoid (right)
Ratio of area 1 0.77 3.46 0.37 0.22
Compute time (s) 2.958 0.017 4.199 0.014 1.520
Ratio of time 176.65 1 250.78 0.84 90.78
@@
Table 3: Comparison of various approaches for Example 6.1. The proposed approach is about two orders of magnitude faster than existing approaches [4, 9], while producing sufficiently accurate approximations.
@@
Figure 1: Robust controllable set for a double integrator system over a horizon of $T=20$ for a ball-shaped $\mathcal{W}$ (left) and an ellipsoidal $\mathcal{W}$ (right) computed via existing approaches (exact [4] and two-stage approach [9]) and the proposed approximations computed using the recursion in Section 5.2. The insets in the left figure show that the proposed approach with ellipsoidal $\mathcal{W}$ (cyan) provides sufficiently accurate inner-approximations of the exact RC set (white).
@@ 6.2 An example with convex polyhedral $\mathcal{X}$
Figure 2: Snapshots of the $100$-step robust controllable sets from various methods for recursion steps $t\in\{0,40,80\}$ (see (43d) in Section 5.1). Our approach computes inner-approximations of the RC set that are similar to those obtained using the exact approach [4] and the two-stage approach [9], with significantly lower computational effort (see Table 4).
@@
and an interval input set $\mathcal{U}=\llbracket-1.5,1.5\rrbracket$, zonotopic disturbance set $\mathcal{W}=(0.01I_{2},[0;0])$, zonotopic goal set $\mathcal{G}=(0.5I_{2},[1.5;0])$, and convex polyhedral, time-invariant state constraints $\mathcal{X}=\{x|[-1,0;2,1]x\leq[2;5]\}$.
Since the state constraints are polyhedral and $A$ is invertible, we generate inner-approximations of the $T$-step RC set using the recursion in Section 5.1. Similarly to Section 6.1, we compare the obtained sets with their exact counterparts computed using MPT3 [4] and the inner-approximations obtained using the two-stage approach [9]. We also compute an outer-approximation of the RC sets using Algorithm 3.
Table 4 shows that the proposed inner-approximating approach with zonotopic $\mathcal{W}$ is over two to three orders of magnitude faster than the existing approaches [4, 9], while providing reasonably accurate inner-approximations that cover about $89\%$ of the area of the exact RC set, as in Section 6.1. The shorter computation time compared to existing approaches is attributed to the optimization-free implementation of Section 5.1, by Theorem 2 and Corollary 1.
Figure 2 shows the $100$-step RC sets computed by various methods at recursion step $t\in\{0,40,80\}$. Unlike in Section 6.1, the proposed inner-approximation of the RC set in this case is contained in the inner-approximation using the two-stage approach [9]. However, the conservativeness of the proposed approach compared to the two-stage approach [9] appears minimal, covering about $89\%$ vs $92\%$ of the area of the exact RC set.
Property Exact $\mathcal{W}$ represented as a zonotope
RC set [4] Proposed $\mathcal{M}^{-}$ Proposed $\mathcal{M}^{+}$ Two-stage [9]
Ratio of area 1 0.89 1.36 0.92
Compute time (s) 110.661 0.090 158.753 10.192
Ratio of time 1235.87 1 1772.96 113.82
Table 4: Comparison of various approaches for Example 6.2. The proposed approach is over two to three orders of magnitude faster than existing approaches [4, 9], while producing sufficiently accurate approximations, even for a long horizon of $T=100$ time steps.
@@ 6.3 Scalability: Chain of damped spring-mass systems
We now demonstrate scalability of the representation complexity for the proposed inner-approximation. Specifically, we compute the RC set of a chain of ${N_{M}}\in\mathbb{N}$ homogenous mass-spring-damper systems, see Figure 3, for a range of chain lengths, i.e., different system dimension, and set recursion lengths, i.e., different horizon $T$. The chain system has the following continuous-time linear time-invariant dynamics,
$\ddot{x}_{N_{M}}$ $=-\frac{2k}{m}x_{N_{M}}+\frac{k}{m}x_{({N_{M}}-1)}-\frac{\mu}{m}\dot{x}_{N_{M}}+w_{N_{M}}+u_{N_{M}},$ (57b)
where $j\in\mathbb{N}_{[2:({N_{M}}-1)]}$. The system (6.3) describes a $(2{N_{M}})$-dimensional system parameterized by the spring constant $k$, the mass $m$, and the friction coefficient $\mu$. Additionally, each spring is actuated by an acceleration input $u_{i}\in\mathcal{U}\subset\mathbb{R}$ and is subject to an acceleration disturbance $w\in\mathcal{W}\subset\mathbb{R}$.
After discretizing (6.3) with sampling time $\Delta T=0.1$ using zero-order hold, we consider the following computations of $T$-step RC sets:
1) $n\in\mathbb{N}_{[4:100]}$ with $T=20$ using Section 5.2,
2) $n\in\mathbb{N}_{[4:100]}$ with $T=40$ using Section 5.2,
@@
Figure 4 shows that the proposed method takes significantly shorter computation time to produce an inner-approximation to the $20$-step RC sets, when compared to existing methods [9, 4]. Specifically, we observe that our approach takes $12.52$ seconds to inner-approximate the $20$-step RC set for a $100$-dimensional system (${N_{M}}=50$). On the other hand, the two-stage approach [9] took $649.74$ seconds (about $10$ minutes) to compute an inner-approximation for the $20$-step RC set for much smaller dimensional system $n=40$ (${N_{M}}=20$). We encountered numerical issues for the exact set computation using MPT3 [4] beyond $n=14$ (${N_{M}}=7$). As expected, the proposed approach took longer to compute the RC sets for an horizon $T=40$ compared to the sets for an horizon $T=20$, but still computed an inner-approximation to the $40$-step RC set for the $100$ dimensional system in $126.53$ seconds (about $2$ minutes). The scalability of the proposed approach compared to existing approaches may be attributed to the optimization-free implementation of Section 5.2, made possible by Theorem 2 and Corollary 1.
We also note that the representation complexity of the proposed inner-approximations grows moderately. For a $10$-dimensional system, an inner-approximating constrained zonotope $\mathcal{K}$ for the $20$-step RC set had a representation complexity of $\mathscr{C}(\mathcal{K})=(10,730,620)$. As noted in Remark 2, one could use exact or approximate reduction techniques (see [7, 6, 25]) to further lower the set representation complexities, if so desired.

### v2

@@ 5.2 Polytopic $\mathcal{X}_{t}$
where (51c) combines (49c) and (49d) into a single step using (7c). Assuming time-invariance, $\mathscr{C}(\mathcal{X})=(L_{X},1)$ and the representation complexity of $\mathcal{K}_{t}$ grows by $({M_{BU}}+L_{X}+n,\mathscr{O}_{{B\mathcal{U}}})$ at each step of the set recursion (see (7c) and Table 2). Consequently, with $\mathscr{C}(\mathcal{G})=({M_{\mathcal{G}}},\mathscr{O}_{{\mathcal{G}}})$, the representation complexity of the inner-approximation of the $T$-step RC set is
$\mathscr{C}(\mathcal{K}^{-})=({M_{\mathcal{G}}}$ $+T({M_{B\mathcal{U}}}+L_{X}+n),\mathscr{O}_{{\mathcal{G}}}+T\mathscr{O}_{{B\mathcal{U}}}).$ (52)
Similarly to (50), $\mathscr{C}(\mathcal{K}^{-})$ in (52) also grows linearly with $T$ and does not depend on the disturbance set $\mathcal{W}$.
@@
We can also use the set recursions discussed in Sec. 5.1 and 5.2 in conjunction with Algo. 5 to obtain an outer-approximation to the $T$-step RC set.
@@ 6 Case studies
We now demonstrate the computational efficiency, scalability, and utility of our approach in several case studies. First, we consider two case studies involving low-dimensional systems to illustrate the advantage of the proposed approach in computing time when compared to existing inner-approximating approaches based on constrained zonotopes [9], and exact approaches based on H-Rep and V-Rep polytopes [24]. Then, we discuss the scalability of the approach by computing the RC set for a chain of mass-spring-damper system, and empirically demonstrate that the proposed approach is numerically stable and can compute the RC sets for high-dimensional systems and long horizons.
We perform the presented computations in a standard computer with Intel CPU i9-12900KF processor (3.2 GHz, 16 cores) and 64 GB RAM, running MATLAB 2022b on Windows. We use YALMIP [28], MOSEK [29], and GUROBI [30] to set up and solve the optimization problems. For two-dimensional plots of constrained zonotopes $\mathcal{C}$, we compute the appropriate H-Rep/V-Rep polytope approximations via support function and vector computations (4) in $100$ equi-spaced directions in $\mathbb{R}^{2}$. We estimate the volume of the sets via grid-based sampling.
6.1 Double integrator example with polytopic $\mathcal{X}$
We consider the computation of RC set for a double integrator system with polytopic state constraints $\mathcal{X}$ and an ellipsoidal disturbance set $\mathcal{W}$. The linear time-invariant system matrices are
@@ 6.1 Double integrator example with polytopic $\mathcal{X}$
with the sampling period $\Delta T=0.1$, the input set $\mathcal{U}=\llbracket-2,2\rrbracket$ that is an interval, the disturbance set $\mathcal{W}=(0.1I_{2},[0;0])$ or $\mathcal{W}=(\text{diag}([0.2,0.04]),[0.1;0.1])\}$ that is a circle or an ellipsoid respectively, and the state constraints $\mathcal{X}=\mathcal{G}=\llbracket-2,2\rrbracket\times\llbracket-3,3\rrbracket$ that are time-invariant, axis-aligned rectangles.
@@
We generate inner-approximations of the $T$-step RC set using the recursion in Sec. 5.2 for $T=20$ with an exact ellipsoidal representation of $\mathcal{W}$ and a zonotopic outer-approximation $\mathcal{W}^{+}=(G_{W},c_{W})\supset\mathcal{W}$ where the Pontryagin difference is inner-approximated using Algo. 3. We also compare the computed sets with the exact sets computed using MPT3 where the vertex-facet enumeration was accomplished using Fourier-Motzkin elimination [24], and the inner-approximations of the RC sets using the two-stage approach with the zonotopic $\mathcal{W}^{+}$ [9]. We also compute an outer-approximation of the RC sets using Algo. 5, as discussed in Rem. 4.
@@
Table 3 shows that the proposed inner-approximation approach with ellipsoidal $\mathcal{W}$ is about two orders of magnitude faster than the exact approach [24] and the $\mathcal{W}^{+}$-based approximation via two-stage approach [9], while providing reasonably accurate inner-approximations (about $97\%$ and $77\%$ of the area of the exact RC set in the first and second case respectively). Our approach with zonotopic $\mathcal{W}^{+}$ is slightly faster than with ellipsoidal $\mathcal{W}$, but it is more conservative due to the zonotopic outer-approximation of the disturbance set. The shorter computation times of the proposed approach are a direct result of Thm. 1 and Prop. 2, since the implementation of the set recursion in Sec. 5.2 can be accomplished in closed-form, i.e., optimization-free. As expected, $({M_{\mathcal{K}_{0}}},\mathscr{O}_{{\mathcal{K}_{0}}})$ of the inner-approximations do not depend on choice of $\mathcal{W}$ (see (52)). The proposed outer-approximation is slower than the inner-approximation, primarily due to the use of linear programming to produce a minimal representation of the convex polyhedron in Step 3 of Algo. 5 to manage the representation complexity. Additionally, ${M_{\mathcal{K}_{0}}}$ for outer-approximation is higher than the inner-approximation due to Step 3 of Algo. 5. For most safe constrained control problems, an inner-approximation of the $T$-step RC set is sufficient.
@@
Fig. 2 shows that the RC sets and their corresponding inner-approximations, associated with a ball-shaped $\mathcal{W}$ (left) and an ellipsoidal $\mathcal{W}$ (right). As expected, the inner-approximations of the RC set constructed using zonotopic $\mathcal{W}^{+}$ are more conservative than their ellipsoid-based counterparts in both cases. On the other hand, the proposed inner-approximations with zonotopic $\mathcal{W}^{+}$ are identical (left) or similar (right) to the inner-approximations produced by the existing two-stage approach [9], while requiring significantly shorter computation time (see Table 3).
Method Area Compute time Complexity $\mathscr{C}$ Area Compute time Complexity $\mathscr{C}$
Ratio Time (s) Ratio ${M_{\mathcal{K}_{0}}}$ $\mathscr{O}_{{\mathcal{K}_{0}}}$ Ratio Time (s) Ratio ${M_{\mathcal{K}_{0}}}$ $\mathscr{O}_{{\mathcal{K}_{0}}}$
$\mathcal{W}$ is a ball (left in Fig. 2) $\mathcal{W}$ is an ellipsoid (right in Fig. 2)
Exact [24] 1 2.044 177.39 N/A 1 2.958 176.65 N/A
Ours $\mathcal{M}^{-}$ 0.97 0.012 1 120 11 0.77 0.017 1 120 11
Ours $\mathcal{M}^{+}$ 1.67 4.064 352.66 322 11 3.46 4.199 250.78 326 11
With a zonotope outer-approximation $\mathcal{W}^{+}\supset\mathcal{W}$ as the disturbance set
Ours $\mathcal{M}^{-}$ 0.70 0.010 0.86 120 11 0.37 0.014 0.84 120 11
2-stage [9] 0.67 1.594 138.29 120 11 0.22 1.520 90.78 120 11
@@
Table 3: Comparison of various approaches for Sec. 6.1. Our (inner-approximation) approach is about two orders of magnitude faster than existing approaches [24, 9], and generates sufficiently accurate approximations.
@@
Figure 2: Robust controllable sets computed using the recursion in Sec. 5.2 for Sec. 6.1 with a ball-shaped $\mathcal{W}$ (left) and an ellipsoidal $\mathcal{W}$ (right). We compare the sets obtained with the proposed approximations to the sets from existing approaches (exact [24] and the two-stage approach [9]). The insets in the left figure show that the proposed approach with ellipsoidal $\mathcal{W}$ (cyan) provides sufficiently accurate inner-approximations of the exact RC set (white).
@@ 6.2 An example with convex polyhedral $\mathcal{X}$
and an interval input set $\mathcal{U}=\llbracket-1.5,1.5\rrbracket$, zonotopic disturbance set $\mathcal{W}=(0.01I_{2},[0;0])$, zonotopic goal set $\mathcal{G}=(0.5I_{2},[1.5;0])$, and convex polyhedral, time-invariant state constraints $\mathcal{X}=\{x\mid[-1,0;2,1]x\leq[2;5]\}$.
Since the state constraints are polyhedral and $A$ is invertible, we generate inner-approximations of the $T$-step RC set using the recursion in Sec. 5.1. Similarly to Sec. 6.1, we compare the obtained sets with their exact counterparts computed using MPT3 [24] and the inner-approximations obtained using the two-stage approach [9]. We also compute an outer-approximation of the RC sets using Algo. 5.
Table 4 shows that the proposed inner-approximating approach with zonotopic $\mathcal{W}$ is over two to three orders of magnitude faster than the existing approaches [24, 9], while providing reasonably accurate inner-approximations that cover about $89\%$ of the area of the exact RC set, as in Sec. 6.1. The shorter computation time compared to existing approaches is attributed to the optimization-free implementation of Sec. 5.1, by Prop. 2 and Corr. 1. As expected, $({M_{\mathcal{K}_{0}}},\mathscr{O}_{{\mathcal{K}_{0}}})$ for our $\mathcal{M}^{-}$ and the two-stage approach are identical [9].
Method Area Compute time Complexity $\mathscr{C}$
Ratio Time (s) Ratio ${M_{\mathcal{K}_{0}}}$ $\mathscr{O}_{{\mathcal{K}_{0}}}$
Exact [24] 1 110.661 1235.87 N/A
Ours $\mathcal{M}^{-}$ 0.89 0.090 1 200 202.0
Ours $\mathcal{M}^{+}$ 1.36 158.753 1772.96 1703 953.5
2-stage [9] 0.92 10.192 113.82 200 202.0
Table 4: Comparison of various approaches for Sec. 6.2. Our (inner-approximation) approach is about two to three orders of magnitude faster than existing approaches [24, 9], and generates accurate approximations even for a long horizon $T$.
Figure 3: Snapshots of the $100$-step robust controllable sets computed using Sec. 5.1 for Sec. 6.2 at recursion steps $t\in\{0,40,80\}$. Our approach computes inner-approximations of the RC set that are similar to those obtained using the exact approach [24] and the two-stage approach [9], with significantly lower computational effort (see Table 4).
Fig. 3 shows the $100$-step RC sets computed by various methods at recursion step $t\in\{0,40,80\}$. Unlike in Sec. 6.1, the proposed inner-approximation of the RC set in this case is contained in the inner-approximation using the two-stage approach [9]. However, the conservativeness of the proposed approach compared to the two-stage approach [9] appears minimal, covering about $89\%$ vs $92\%$ of the area of the exact RC set.
@@ 6.3 Scalability: Chain of damped spring-mass systems
We now demonstrate scalability of the representation complexity for the proposed inner-approximation. Specifically, we compute the RC set of a chain of ${N_{M}}\in\mathbb{N}$ homogenous mass-spring-damper systems, see Fig. 4, for a range of chain lengths, i.e., different system dimension, and set recursion lengths, i.e., different horizon $T$. The chain system has the following continuous-time linear time-invariant dynamics,
$\ddot{x}_{N_{M}}$ $=-\frac{2k}{m}x_{N_{M}}+\frac{k}{m}x_{({N_{M}}-1)}-\frac{\mu}{m}\dot{x}_{N_{M}}$
$\qquad\qquad\hskip 8.00003pt+w_{N_{M}}+u_{N_{M}},$ (63b)
where $j\in\mathbb{N}_{[2:({N_{M}}-1)]}$. Here, (63) describes a $(2{N_{M}})$-dimensional system parameterized by the spring constant $k$, the mass $m$, and the friction coefficient $\mu$. Each spring is actuated by an acceleration input $u_{i}\in\mathcal{U}\subset\mathbb{R}$ subject to an acceleration disturbance $w\in\mathcal{W}\subset\mathbb{R}$.
@@
After discretizing (63) with sampling time $\Delta T=0.1$ using zero-order hold, we consider the following computations of $T$-step RC sets:
1) $n\in\mathbb{N}_{[4:100]}$ with $T=20$ using Sec. 5.2,
2) $n\in\mathbb{N}_{[4:100]}$ with $T=40$ using Sec. 5.2,
@@
Fig. 5 shows that the proposed method takes significantly shorter computation time to produce an inner-approximation to the $20$-step RC sets, when compared to existing methods [9, 24]. Specifically, we observe that our approach takes $12.52$ seconds to inner-approximate the $20$-step RC set for a $100$-dimensional system (${N_{M}}=50$). On the other hand, the two-stage approach [9] took $649.74$ seconds (about $10$ minutes) to compute an inner-approximation for the $20$-step RC set for much smaller dimensional system $n=40$ (${N_{M}}=20$). We encountered numerical issues for the exact set computation using MPT3 [24] beyond $n=14$ (${N_{M}}=7$). As expected, the proposed approach took longer to compute the RC sets for an horizon $T=40$ compared to the sets for an horizon $T=20$, but still computed an inner-approximation to the $40$-step RC set for the $100$ dimensional system in $126.53$ seconds (about $2$ minutes). The scalability of the proposed approach compared to existing approaches may be attributed to the optimization-free implementation of Sec. 5.2, made possible by Prop. 2 and Corr. 1.
We oberved a moderate growth in the representation complexity of the proposed inner-approximations. For a $10$-dimensional system, an inner-approximating constrained zonotope $\mathcal{K}_{0}$ for the $20$-step RC set had a representation complexity of $\mathscr{C}(\mathcal{K}_{0})=(620,11)$ with $n=10,{N_{\mathcal{K}_{0}}}=730,{M_{\mathcal{K}_{0}}}=620$. As noted in Rem. 5, various reduction techniques may be used to further lower the set representation complexities, if so desired.